Determine whether the equation defines a function with independent variable x. If it does, find the domain. If it does not, find a value of x to which there corresponds more than one value of y.
x|y| = x + 9

Found 2 solutions by user_dude2008, solver91311:
Answer by user_dude2008(1862)

x|y| = x + 9

|y| = (x + 9)/x

y = (x + 9)/x .... or y = -(x+9)/x

the equation is not a function of x

x=2 ----> y=(2+9)/2=11/2 or x=2 ---> y = -(2+9)/2 = -11/2

Answer by solver91311(24713)







Using the definition of absolute value:



We can write:

OR

Let , then



OR



Therefore, NOT a function


John